# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v1

@@ Abstract
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that inject insulin automatically. However, insulin is a dangerous hormone, where too much insulin can kill people in a matter of hours and too little insulin can kill people in a matter of days.
In this paper, we take on the challenge of building a new trustworthy automated insulin delivery system, called GlucOS. In our design, we apply separation principles to keep our implementation simple, we use formal methods to prove correct the most critical parts of the system, and we design novel security mechanisms and policies to withstand malicious components and attacks on the system.
We report on real world use for one individual for 6 months using GlucOS. Our data shows that for this individual, our ML-based algorithm runs safely and manages their T1D effectively. We also run our system on 21 virtual humans using simulations and show that our security and safety mechanisms enable ML to improve their core T1D measures of metabolic health by 4.3% on average. Finally, we show that our security and safety mechanisms maintain recommended levels of control over T1D even in the face of active attacks that would have otherwise led to death.
GlucOS is open source and our code is available on GitHub.
@@ 1 Introduction
Type 1 diabetes (T1D) is an autoimmune disease in which an individual’s pancreas stops producing insulin. People living with T1D must therefore take synthetic insulin and play the role of their pancreas. However, replicating the pancreas’s native insulin release patterns is extremely difficult, and optimal T1D self-management often requires hourly adjustments in insulin dosing both in response to physiological changes and in prediction of upcoming carbohydrate intake and energy expenditure (e.g., exercise, stress) [51, 56]. Even when following this labor-intensive process of T1D self-management, affected individuals remain at risk for acute episodes of hypoglycemia (low blood sugar, which can cause loss of consciousness, seizure, or death) or ketoacidosis (which carries an associated risk for multi-organ injury), as well as for long term vascular complications (e.g., kidney damage, nerve damage, blindness, heart attack, stroke) that are a result of chronic hyperglycemia (high blood sugar) [46]. 8.4 million people are living with T1D worldwide [25].
(a) Devices for automated insulin delivery. This figure is from jdrf.org.
@@
Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and apps.
The last two decades have witnessed the development of therapeutic technology to assist T1D management, including continuous glucose monitoring (CGM) devices and insulin pumps. Most recently, the integration of CGM sensors and insulin pumps into “artificial pancreas” or automated insulin delivery systems has enabled a simultaneous improvement in health outcomes and reduction in the burden of care for T1D-affected individuals [38].
However, existing automated insulin delivery systems are complex, and there is no way to assure beyond any doubt that they are implemented correctly. For example, techniques from formal methods have not been applied to verify automated insulin delivery systems. With little reason to trust the systems and their core algorithms, users and developers are also reluctant to incorporate experimental algorithms into the prediction loop, such as advanced ML-driven insulin delivery. Researchers have shown that ML can improve automated insulin delivery systems [64, 33, 63, 59, 65, 58], yet the lack of security and safety for these ML predictions keeps people from using it. If ML is going to be part of medical decisions, we need the security community to lead the way on how to do this safely.
In this paper, we argue for applying traditional security and operating systems design principles to make automated insulin delivery systems more trustworthy. This includes going beyond just patient privacy or medical device security, to systems security for automated insulin delivery systems to ensure that they are implemented correctly and that the incorporation of ML does not jeopardize their safety.
@@
Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference. However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
Designing and implementing trustworthy automated insulin delivery systems brings with it three main challenges. (1) How can automated insulin delivery design decisions improve rather than worsen the security posture of such software for end users? (2) How do we apply formal methods to a complex, distributed system that includes a human? (3) How can we cope with ML mispredictions, including those from malicious ML?
In this paper, we take on these challenges and design and implement a new trustworthy automated insulin delivery system, called GlucOS. Our design for GlucOS strives to achieve the following four primary goals: to use sound security principles to separate security and safety concerns into critical and non-critical components; to keep our implementation as simple as possible; to prove correct the most critical components; and to enable any predictive algorithms, including deep neural network ML, to control the insulin pump securely and safely.
GlucOS improves on the current state-of-the-art in automated insulin delivery systems by, for the first time, enabling deep neural network ML to control the insulin pump securely and safely. Our key insight for security is that over a long enough period of time, all correct algorithms will dose the same amount of insulin. For example, on a given day, a person will need to dose a fixed amount of insulin to absorb all of the glucose from the food they eat. Since synthetic insulin action is slow, injecting insulin early by predicting future metabolic states is critical for long-term health. GlucOS’s support for modern ML-based algorithms provides users with this predictive power to dose insulin. For security, GlucOS pairs a predictive ML model with a conservative and safe model to provide the ML with enough flexibility to control the insulin pump proactively while staying within the bounds dictated by the safe model.
Our architecture and system are a stark contrast to current systems. Current systems are complex, their architecture provides little or no isolation boundaries, they are unable to withstand attacks on their system, and they are unable to use deep neural network ML securely and safely. Given that mistakes with insulin delivery can kill people, our goal is to have a fully-featured automated insulin delivery system with security and safety as the primary design considerations. Even if parts of the system are malicious, GlucOS continues to operate and ensure that the individual using the system can manage their T1D safely.
We report our experiences running GlucOS on one individual for six months to manage their T1D. Running our software on a real human forces us to design a practical and real system. Our evaluation using virtual humans in a simulator shows that our security and safety mechanisms generalize beyond the specific individual.
Our novel contributions include:
  • The clean-slate design and implementation of a simple yet effective automated insulin delivery system.
@@
  • A case study describing our experiences from a real-world deployment, and results from simulation to show that our techniques generalize.
We give top priority to the health, safety, and ethics of the human using GlucOS to manage their T1D. Section 11.1 summarizes our safety protocol and Appendix A provides a detailed description.
2 Background
Type 1 Diabetes. T1D is an autoimmune condition in which the immune system attacks and destroys the insulin-making cells in the pancreas. Insulin is a hormone that enables the absorption of glucose (sugar) from the bloodstream into the brain, the muscles, and other organs and tissues where it is used as an energy source. When people eat, the digestive system converts food into glucose, which makes its way into the bloodstream, but it cannot then be used by the body without the action of insulin.
Since people with T1D are unable to produce their own insulin, they must inject synthetic insulin to replace this key function, and the appropriate dosing of this insulin is both complicated and labor-intensive. For example, insulin doses must be closely correlated with food intake, but the absorption of exogenous insulin is slower than food digestion, so people with T1D must precisely predict their future food intake in order to inject the right amount of insulin prior to each meal. Once injected, typical insulin formulations will affect glucose metabolism for up to three hours, and during that time will interact with other factors like exercise, sleep, stress, caffeine, and alcohol. However, the exact effects of these factors on an individual’s glucose levels are even harder to predict than those of food alone, resulting in the need for continuous adjustments by the individual in order to maintain a balanced glycemic state. One solution that has emerged to address the complications of daily T1D management is automated insulin delivery systems, which connect a CGM sensor and insulin pump to enable automatic, individualized adjustments in insulin delivery based on glucose trends [14, 18, 40, 48, 17, 7, 4].
The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes. Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness and seizure, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage. The main physiological concept in this paper is a goal of maintaining glucose levels that are in the target range the majority of the time.
Automated insulin delivery systems. Automated insulin delivery systems connect a CGM device – which measures interstitial glucose concentration every 1-5 minutes – with an insulin pump, which delivers a continuous subcutaneous insulin infusion. Using a variety of algorithms, these systems vary insulin delivery in response to CGM data, reducing the amount of human input and cognitive load required to achieve glycemic stability (Figure 1). Several commercial [17, 7, 4] and open source [40, 48] automated insulin delivery systems exist today.
@@ 3 Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery. In this section, we first provide an overview of GlucOS’s system architecture to describe the key components involved in automated insulin delivery. While we provide an overview of all components, we specifically focus on those components that are part of our trusted computing base. We then define our threat model in terms of our assumptions and goals with GlucOS.
@@ 3.1 Architecture
From a system design perspective, our primary goals are to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas [10, 30] applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers [26, 61, 57, 53], which also apply OS separation principles at the application layer. By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple. Also, we want flexibility so that people can run whatever closed-loop algorithms they want to and use whatever pump hardware they want – we focus on providing mechanisms, policies, and systems for security and safety.
@@
The BioKernel runs the closed-loop algorithm every five minutes as new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model. The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
@@
For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a reactive safe physiological model. For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump. These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
@@ 3.2 Threat Model
We consider attacks from malicious predictive ML algorithms as well as malicious pump drivers. We assume that malicious ML can intentionally miscalculate the amount of insulin to inject and malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
We also consider real-world safety concerns as part of our threat model. ML mispredictions are part of these safety concerns since even well-designed ML will sometimes make poor decisions. Safety concerns also include pump communication errors emerging from bugs in the pump driver, pump mechanical failures [55], and pump site skin inflammation or infection [36], all of which result in a mismatch between the amount of insulin calculated in closed loop and the actual amount of insulin active in the human.
Our goal with GlucOS is to not merely detect attacks and safety violations. Our goal is to build a secure system for automated insulin delivery that can withstand active attacks and real-world safety concerns.
We assume that CGM sensors are correct and operate within their error bounds. If the CGM sensors stop producing measurements, our system will be unable to handle it. We also assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware (such as adversarial attacks against predictive ML) in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
@@ 4 Algorithmic security: Withstanding malicious ML
None of the current automated insulin delivery systems use deep neural network ML. Although they do use ML in general, they use statistical methods, like linear regression or model predictive control [19, 60, 42, 52, 48], to learn physiological parameters. The reason for using this more traditional form of ML is sound: it provides determinism and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [64, 33, 63, 59, 65, 58]. Despite this promise, none of the current automated insulin delivery systems use more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could be lethal if it delivers an inappropriate insulin dose [20, 13].
Modern ML models make mistakes. High-profile examples include Google Photos classifying black people as gorillas [2], LinkedIn recruiting ML learning gender bias [5], Microsoft’s chatbot, Tay, trained on Twitter data (what could go wrong) becoming racist [3], and ChatGPT large language model hallucinations [1]. In all of these examples, professional ML engineers with plenty of resources produced well-designed ML systems that hit a gray area in their training data or decision space and made bad decisions.
All automated insulin delivery systems that use ML need to have a mechanism to protect against malicious ML or benign ML that makes inevitable bad decisions. In this section, we discuss the design and implementation of GlucOS’s algorithmic security and safety mechanisms.
@@ 4.1 Mechanism fundamentals
Our key insight that underpins our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a long enough period of time. However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 10). Thus, our algorithmic security mechanism pairs a predictive model that anticipates future metabolic states and doses insulin accordingly with a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time, and we use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get both the safety of a reactive safe model that operates only on facts (Section 4.2) and the tight control we expect from an ML-based predictive model (Section 4.3).
@@
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model. These models independently produce pump commands, called a temporary basal command. The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration. The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model. If the predictive ML model is still within its insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump. If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model. For more details about how this algorithm works, please see our implementation in GitHub (link omitted).
4.2 Reactive safe model
Since we use our reactive safe model to anchor our security policy, and it runs within our trusted computing base, our goals for this model are to have it be safe, easy to understand, and have a simple implementation.
To make this model safe, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated. From this calculation, we know how much insulin to inject or withhold.

### v2

@@ Abstract
We present GlucOS, a novel system for trustworthy automated insulin delivery. Fundamentally, this paper is about a system we designed, implemented, and deployed on real humans and the lessons learned from our experiences.
GlucOS combines algorithmic security, driver security, and end-to-end verification to protect against malicious ML models, vulnerable pump drivers, and drastic changes in human physiology. We use formal methods to prove correctness of critical components and incorporate humans as part of our defensive strategy. Our evaluation includes both a real-world deployment with seven individuals and results from simulation to show that our techniques generalize. Our results show that GlucOS maintains safety and improves glucose control even under attack conditions. This work demonstrates the potential for secure, personalized, automated healthcare systems.
@@ 1 Introduction
Type 1 Diabetes (T1D) is a metabolic disorder where an individual’s pancreas stops producing insulin. To compensate, they inject synthetic insulin. Computer systems, called automated insulin delivery systems, exist that use subcutaneous sensors to continuously monitor glucose concentrations (glucose is the body’s primary energy source) and inject insulin automatically using an insulin pump device (Figure 1). However, insulin is a dangerous hormone where too much insulin can kill people in a matter of hours [22] and too little insulin can kill people in a matter of days [13]. For automated insulin delivery systems, the key challenge is to maintain a balance: provide enough insulin to prevent dangerously high glucose levels while avoiding excessive insulin that could lead to life-threateningly low glucose levels. In this context, security means ensuring integrity for insulin dosing against malicious attacks or inadvertant errors. Several commercial [17, 5, 3] and open-source [42, 54] automated insulin delivery systems exist today, but none of them consider security as a primary design constraint.
(a) Devices for automated insulin delivery. This figure is from breakthrought1d.org.
@@
Figure 1: Overview of automated insulin delivery devices, closed-loop architecture, and a screenshot of GlucOS.
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations [71, 72]. We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints. Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system. Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components. Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. To address these challenges, we need practical systems that move beyond simulation-based designs and tackle the challenges of real-world systems.
Current open-source automated insulin delivery systems are complex. For example, the OpenAPS [54] core function for calculating insulin doses consists of 1192 lines of Javascript code, 63 input and configuration parameters, and 90 branch statements. All of this logic influences the ultimate dosing decision, making reasoning about its correctness, both manually and automatically, challenging. Automated insulin delivery systems that use the OpenAPS algorithm embed it into native apps using a WebView for Javascript interpretation, adding to the overall complexity and increasing the attack surface. In contrast, our implementation of this same function is 25 lines of native and formally verified Swift code. None of the open-source automated insulin delivery systems use formal methods, can withstand attacks on components, or have mechanisms or policies for security.
In this paper, we take on the challenge of building the first trustworthy automated insulin delivery system, called GlucOS. Our design includes: (1) an architecture where we apply separation principles for isolated and simple components, (2) a novel security mechanism and policy to enable secure ML-driven insulin delivery, (3) the ability to withstand attacks on insulin pump driver software, and (4) the application of formal methods to prove correct our insulin delivery path, the most sensitive part of the system. Despite these efforts, we identify scenarios in automated insulin delivery that require users to take action for security. Thus, we incorporate humans as part of our defensive strategy, providing them with agency and trust to take the required, corrective actions. Our approach is distinguished by its holistic consideration of the entire problem of automated insulin delivery. We provide clear security boundaries, mechanisms and policies to mitigate attacks on vulnerable components, and formal methods to provide assurances that our implementation is correct.
@@
Automated insulin delivery systems are different from traditional applications that the security community focuses on, but they are important. Research on operating systems or web browsers can list scores of CVEs from current systems and show how their system improves security by eliminating the impact of some or all of these vulnerabilities by virtue of their design. Automated insulin delivery systems do not have a long list of CVEs that we can reference, although there are a few [46, 47, 48, 49]. However, our position is that we need to make sure that these systems never get CVEs, because if they do, security vulnerabilities can kill people. Thus, we focus on building a new system, using security first principles and formal methods, to avoid security vulnerabilities and mitigate their damage from the start.
Our main contribution lies in a new end-to-end system that we built and the lessons learned from our experience designing, implementing, and deploying GlucOS to humans to manage their T1D. Our novel contributions include:
@@
  • A biological invariant that combines end-to-end physiological security checks with formal methods.
  • A case study describing our experiences from a real-world deployment with seven individuals, and results from simulation to show that our techniques generalize.
  • A human-centric design that considers the user as an integral part of the system, influencing both the overall architecture and including them in our defensive strategy.
We give top priority to the health, safety, and ethics of the humans using GlucOS to manage their T1D. Section 9.1 summarizes our safety protocol and Appendix A provides a detailed description.
There are 8.4 million people living with T1D worldwide [30], underscoring the need for providing trustworthy systems for automated insulin delivery. Our work, carried out with rigorous medical and ethical oversight, is a first step in addressing this need.
@@ 2 Overview
This paper describes our design for GlucOS, a system for trustworthy automated insulin delivery; we have two primary goals. First, we want the software to be simple and correct. Second, we want to identify the most vulnerable parts of the system and be able to prevent attacks on these components or withstand successful attacks. This section describes our overall architecture for GlucOS that strives to achieve these goals, and our threat model.
2.1 Locking down the insulin delivery path
We focus our system architecture on providing security mechanisms and policies on the insulin delivery path. The insulin delivery path includes models for calculating the amount of insulin still active in the individual (insulin can take up to six hours to fully absorb), closed-loop algorithms for calculating insulin dosing, deep neural network ML to predict future metabolic states, and drivers for programming the insulin pump. Together, these components are both the most complex, and the most sensitive from a security perspective because they are responsible for insulin delivery.
@@ 2.2 Architecture
From a system design perspective, we strive to keep our implementation simple and enable our use of formal methods to prove correct the most critical parts. For our software system architecture, we use separation principles from the OS and microkernel areas [8, 33] applied to the application layer to push complexity away from the most critical parts of the system. This architecture is similar to secure web browsers [31, 67, 63, 59], which also apply OS separation principles at the application layer. By decomposing the overall automated insulin delivery system, we strive to keep our trusted computing base simple.
@@
The BioKernel runs the closed-loop algorithm every five minutes as the new CGM readings come in. Once the BioKernel receives new CGM readings, it collects the current metabolic state of the individual and sends it to a predictive ML model. The predictive ML model calculates the amount of insulin to inject based on its assessment of the individual’s current metabolic state and predictions of future metabolic states. The BioKernel then issues commands to the pump to deliver the calculated amount of insulin. A pump driver converts these commands into low-level I/O to program the pump and adjust insulin delivery.
@@
For predictive ML, our reference monitor enforces algorithmic security, where we bound the amount of insulin calculated by ML to stay within dynamic safety bounds computed by a simple reactive safe model. For the pump driver, we interpose on the I/O path to ensure that only the pump commands issued by the BioKernel are programmed into the pump. These checks constitute our pump invariants, which, when violated switch the pump to operate in a safe mode.
@@ 2.3 Threat model
The most serious health risk for people who use synthetic insulin is low glucose, or hypoglycemia. Hypoglycemia results from an overdose of insulin. Once an automated insulin delivery system has delivered insulin, there is no way to remove it from the body. If untreated, severe hypoglycemia can lead to loss of consciousness, seizure, or death in a matter of hours. Thus, our primary security objective is to reduce or eliminate hypoglycemia that results from an attack on the system. Quantitatively, our goal is to maintain the American Diabetes Association’s recommended time in hypoglycemia of less than 4% [12].
As our design philosophy centers around humans, providing flexibility for people to use whatever ML they want by loading personalized models or using any pump hardware they want is important. This flexibility helps guide our focus on providing novel security mechanisms and policies for these components. Our goal is to make sure that they stay safe even if the model they pick is malicious or the pump driver they choose has security vulnerabilities. Attackers may target automated insulin delivery systems because they can directly harm the human using the system.
We assume that the Swift type system is correct and that the underlying operating system, the CGM driver and hardware, and the pump hardware we use are free of attacks and their device communications maintain integrity. Thus, we do not consider attacks emerging from tampered CGM sensors or pump hardware in our threat model. These are important but complementary problems to our work, which focuses on the system architecture and algorithms for automated insulin delivery.
We consider attacks from malicious predictive ML algorithms. These attacks can come from trojan attack logic embedded directly in a model, or from an adversarial attack on predictive ML. For adversarial attacks, although we assume that the CGM and insulin pump remain attack free, an adversary can manipulate these entities by serving an individual who is using our system with calculated amounts of carbohydrates in their food (e.g., at a restaurant), thus influencing the resulting CGM readings and corresponding insulin commands. In our threat model, malicious predictive ML can issue arbitrary insulin doses.
We also consider malicious pump drivers. In our threat model, malicious pump drivers can issue arbitrary pump I/O operations to add, drop, or modify commands sent to the pump.
Successful attacks on predictive ML or pump drivers, if unchecked, can be devastating because they directly affect insulin delivery and can be used by an attacker to induce potentially life-threatening hypoglycemia.
@@ 3 Algorithmic security: Withstanding malicious ML
Currently, no automated insulin delivery systems use deep neural network ML. Some current systems use statistical methods, like linear regression or model predictive control [21, 66, 43, 58, 54]. The reason for using this more traditional form of ML is sound. It provides consistent, explainable results and has a physiological basis that people can use to vet its decisions. However, all of the recent advances in ML, like image classifiers and large language models, have come from deep neural networks, which researchers have shown also work well for predicting metabolic states for use in automated insulin delivery systems [73, 35, 70, 65, 74, 64]. Despite this promise, none of the current automated insulin delivery systems incorporate more advanced deep neural networks due to the black-box nature of deep neural networks and the potentially dire consequences of mispredictions – an automated insulin delivery system’s unchecked “hallucination” [1] could have lethal consequences if it delivers an inappropriate insulin dose [22, 13].
In this section, we discuss the design and implementation of GlucOS’s algorithmic security mechanisms to enable predictive models, including deep neural network ML, to run safely in automated insulin delivery systems.
@@ 3.1 Mechanism fundamentals
Our key novel insight underpinning our security mechanism design is that all correct insulin delivery algorithms will deliver the same amount of insulin over a sufficiently long period. People inject insulin to facilitate glucose absorption after eating. Digestion converts food to glucose, and insulin facilitates the absorption of glucose to tissue (e.g., the brain) where the body can use it as an energy source. Since insulin dosing is driven by food consumption, all correct algorithms should dose the same amount of insulin to cover glucose from digestion.
However, the timing of this insulin delivery matters for maintaining balanced glucose levels (Section 8). Insulin activation is a slow process, with injected insulin taking up to six hours to fully activate. To meet this challenge, our algorithmic security mechanism pairs two models. Our first model is a predictive model that anticipates future metabolic states and doses insulin proactively. Our second model is a reactive safe model that measures the current metabolic state and doses insulin based only on what it can measure currently. This pairing allows us to balance proactive insulin dosing for faster responses with a more conservative approach for safety.
Architecturally, in a novel approach for automated insulin delivery systems, the predictive model runs outside of our trusted computing base while the reactive safe model is part of the trusted computing base. This separation enables individuals to use whatever predictive models they want, while maintaining security. It also defends against malicious predictive models, as all insulin delivery decisions are ultimately vetted by the reactive safe model. This design also enables people to update their predictive models without violating core safety guarentees.
Conceptually, the predictive model is the primary model in our system and should control the pump most of the time. We use the reactive safe model for accounting so that we can track how far the predictive model deviates from a known-to-be-safe baseline. Then, we bound the size of this deviation. By combining these two models, we strive to get beneficial properties from both. Our approach combines the safety of a reactive safe model and the performance from a predictive model.
Our security mechanism draws inspiration from previous work on combining simple and complex models for safety [61, 51]. However, our approach extends these ideas by providing a novel security mechanism specifically designed for insulin delivery models. Our key contributions lie in our insight into the security properties of insulin delivery algorithms and our principles for how to apply our mechanism to automated insulin delivery systems (Section 3.2).
@@
Figure 3 shows how our algorithmic security mechanism works. The whole process starts when a new CGM reading arrives. Once the BioKernel receives a new CGM reading, it collects a recent history of CGM readings and insulin dosing and sends these as inputs to both the reactive safe model and the predictive ML model. These models independently produce pump commands, called a temporary basal command. The temporary basal command sets the insulin delivery rate for the insulin pump for a 30-minute duration. The BioKernel will collect both temporary basal commands and look back at the differences between the reactive safe model and predictive ML model for the past three hours to see how much insulin delivery it should attribute to the predictive ML model. If the predictive ML model is still within its safe insulin delivery limits, it uses the predictive ML model’s temporary basal command to program the pump. If it has exhausted its limits, we clamp it to fit within the insulin bounds we set relative to the reactive safe model.
To set bounds, we base our calculations on human physiology. People produce a background level of glucose for energy. People living with T1D need to inject insulin to facilitate the absorption of this background glucose.
Our method for setting bounds allows the predictive model to use up to three hours worth of background insulin in advance. If the prediction is correct and glucose levels rise (e.g., from eating), the proactive insulin will manage it. If the prediction is wrong and no extra glucose appears, the body’s ongoing glucose production will help use the extra insulin over the next three hours.
In the worst case, an incorrect prediction might lead to low glucose. However, with our mechanism in place, we limit the impact by our three-hour bound and offset by the body’s continuous glucose production.
3.2 Model principles
Our contribution is not in the specific algorithms we use, but rather in the insight around insulin delivery being equal for all correct algorithms and how we use this insight for security. In this section, we outline the algorithms we use in our current implementation and define the principles for how to apply security to any predictive insulin dosing algorithms.
Algorithmically, we use a PID controller [25] from feedback control for our reactive safe model, and we use a deep neural-network from the literature for meal predictions for our predictive ML model [52]. Although we use this model in our current implementation, our system supports any predictive model. PID controllers are used in automated insulin delivery systems [68, 38] and research predictive ML models for insulin delivery have been well covered in the literature [73, 35, 70, 65, 74, 64].
The design of the reactive safe model is crucial for algorithmic security, as it serves as the anchor for the security mechanism. The ideal reactive safe model should be safe, easy to understand, and have a formally verified implementation (Section 7).
To ensure the safety of this model, we base its calculations solely on facts observed automatically from the CGM and the insulin pump. Given these high-quality sources of data, and a simple insulin absorption physiological model, we can calculate how much insulin an individual needs and how much insulin they have injected but have not yet activated. From this calculation, we know how much insulin to inject or withhold.